I need help finding the formula and how to work this problem. The members of a traveling club plan to share equally the cost of $150,000 motor homw. If they can find 10 more people to join and share the cost, then the cost per person will decrease by 1,250. How many members are ther initially in the club? I know c=150,000/x or c=150,000/x+10 I think. Thanks
Answer by stanbon(75887)

I need help finding the formula and how to work this problem. The members of a traveling club plan to share equally the cost of $150,000 motor home. If they can find 10 more people to join and share the cost, then the cost per person will decrease by 1,250. How many members are ther initially in the club? I know c=150,000/x or c=150,000/(x+10)
-----------------------------------
Right. And the difference between them is $1,250.
So you get this equation:
old average - new average = 1250
(150,000/x) - (150000/(x+10)) = 1250
Divide thru by 1250 to get:
120/x - 120/(x+10) = 1
Multiply thru by x(x+10) to get:
120(x+10) - 120x = x(x+10)
1200 = x^2+1x
x^2+10x-1200 = 0
(x+40)(x-30) = 0
Positive solution:
x = 30 (original number of members)
==================================
Cheers,
Stan H.